# Vibrationally Assisted Exciton Transfer in Open Quantum Systems with Long-Range Interactions

### v1

@@ Abstract
The interplay between coherence and system-environment interactions is at the basis of a wide range of phenomena, from quantum information processing to charge and energy transfer in molecular systems, biomolecules, and photochemical materials. In this work, we use a Frenkel-exciton model with long-range interacting qubits coupled to a damped collective bosonic mode to investigate vibrationally assisted transfer processes in donor-acceptor systems featuring internal substructures analogous to light-harvesting complexes. We find that certain delocalized excitonic states maximize the transfer rate and that the entanglement is preserved during the dissipative transfer over a wide range of parameters. We investigate the transfer reduction by static disorder and by finite temperature and study how transfer efficiency scales as a function of the number of dimerized monomers and the component number of each monomer, finding which excitonic states lead to optimal transfer. Finally, we provide a realistic experimental setting to realize this model in analog trapped-ion quantum simulators. Analog quantum simulation of systems comprising many and increasingly complex monomers could offer valuable insights into the design of light-harvesting materials, particularly in the non-perturbative intermediate parameter regime examined in this study, where classical simulation methods are resource-intensive.
@@ I Introduction
CT and ET in the presence of a complex environment can be generally described by Frenkel-exciton models [30, 31, 32, 33] that employ discrete Hilbert spaces to encode the electronic states along with harmonic oscillators to describe molecular and solvent vibrations. Recently, it has been shown that both electronic and vibrational degrees of freedom can be directly mapped onto state-of-the-art analog quantum simulators based on trapped atomic ions [34] and superconducting qubits [35]. These platforms allow for precise control of the coherent part of the evolution [36, 37, 38, 39, 40, 41, 42] and, in some cases, also provide turning knobs to tailor the system-bath coupling implementing reservoir engineering [43, 44, 45, 46].
In this work, we theoretically and numerically investigate a minimal Frenkel-exciton model, where the donor and acceptor have an internal substructure [47] similar to light-harvesting complexes in which different monomers feature strongly coupled molecular sites. Here, the molecular sites are encoded in collections of two-level systems, among which the excitation can be shared. The two-level systems are coupled to a damped collective bosonic mode, where the damping arises from the presence of an Ohmic bath.
Figure 1: Model and Trapped-Ion implementation: (a) A chain of 4 qubit ions and 2 coolants can realize the Frenkel-exciton model (see main text). The red (blue) ions on the sides represent donor and acceptor monomers, respectively, and the information in each ion is encoded in a qubit. The two cyan ions in the middle represent the cooling ions used for the sympathetic cooling of the ion chain motion. (b) Pictorial representation of the spectrum of the total Hamiltonian in the symmetric case, described by Eqs. (10) and (12). Donor (red) and acceptor (blue) surfaces are shown as a function of the normalized reaction coordinate $y/y_{0}=1/2(a^{\dagger}+a)$ with their respective non-interacting harmonic wavefunctions. The effect of the bath is illustrated by the cyan wavy arrows that represent the relaxation with rate $\gamma$. The qubits in the donor/acceptor sites are strongly coupled and can be described in terms of local triplet and singlet states $|T_{D,A}\rangle,|S_{D,A}\rangle$ separated by an energy $2J$. These states are, in turn, coupled by inter-monomer couplings $J_{AB}$, where $AB=TT,ST,SS$ (see main text). The inter-monomer couplings $J_{AB}$ induce anti-crossings between the two surfaces allowing the donor to acceptor transfer.
@@
In addition to its fundamental theoretical interest, this model is also motivated by being experimentally realizable on an analog trapped-ion simulator, where the electronic degrees of freedom can be encoded in the internal atomic states, and the molecular vibrations in the collective vibrations of the ions around their equilibrium positions (see Section VI). Finally, in this work, we provide experimentally accessible parameters to realize this model in a trapped-ion simulator using a chain of qubit and coolant ions [44]. This paper is organized as follows. In section II, we introduce the model for excitation transfer. Section III focuses on the transfer rate of an initially delocalized excitation in a dimer, where each monomer is realized by two ions, both in the perturbative and non-perturbative regimes. Additionally, we investigate the effect of static disorder on the transfer rate. In section IV, we investigate the effect of temperature on transfer rates. In section V, we discuss how this model can be extended to longer chains with more monomers and with monomers composed of more than two qubits. In section VI, we propose an experimental implementation of the model in a trapped-ion simulator. Finally, in section VII, we summarize our findings and provide an outlook for future studies.
@@ II Model
We consider a one-dimensional system of $N$ spin-$1/2$ qubit ions interleaved with $M$ coolant ions, which are used to perform reservoir engineering on the shared bosonic mode [44], as shown in Fig 1(a). The qubit ions interact via long-range spin exchange interactions $J_{ij}$ [49] and are subject to local energy shifts $\epsilon_{i}$. At the same time, the two-level systems are coupled to the collective bosonic mode defined by creation (annihilation) operators $a^{\dagger}(a)$ with frequency $\omega$, resulting in the total Hamiltonian:
@@
The Hamiltonian $H_{\rm el}$ encodes the electronic degrees of freedom with the energy landscape that determines the roles of the molecular sites and their mutual electronic couplings. We assume a power-law decay for the electronic couplings $J_{ij}=J/d_{ij}^{p}$, that can be natively realized in trapped-ion systems with $J>0$ [49]. Here, $d_{ij}$ is the distance between the qubit ions, and $p$ is the power-law exponent. Systems in $\mathcal{D}$ dimensions with $p\leq 2\mathcal{D}+1$ have generated a lot of interest recently because deriving tight bounds on the ultimate speed of quantum information transfer is challenging in this regime [50, 51, 52, 53, 54]. Here, we consider $p=1$ for most results, which is well within this region and it is a significantly longer range than the dipole-dipole coupling $(p=3)$ typically assumed in Frenkel-exciton models. The Hamiltonian $H_{\rm v}$ describes a collective vibrational mode of motion that can be encoded in a normal mode of the ion chain, defining a reaction coordinate $y/y_{0}=1/2(a^{\dagger}+a)$, with $y_{0}=\sqrt{\hbar/2m\omega}$, and $m$ being the ion mass [55]. Finally, $H_{\rm el-v}$ represents the interaction between electron and vibration with site-dependent couplings $g_{i}$.
@@ III Two-Monomer model
Let us consider a model with only two monomers, which has been used as a minimal model of charge transfer in photosynthetic complexes [59, 23, 22]. In this case, each monomer is composed of two sites and separated by two coolant ions (see section VI.2), as shown in Fig. 1. In this case, we can use the following notation:
@@
Figure 2: Excitation transfer: Transfer rate $k_{T}$ in the perturbative ((a) and (b)) and non-perturbative regimes ((c) and (d)). (a) Transfer rate $k_{T}$ as a function of $\epsilon$ in the perturbative regime ($J,\gamma<<\omega$) for the different initial states $|T_{D}\rangle$ (T), $|S_{A}\rangle$ (S), and $|2\rangle$ (P). The black solid line represents the results predicted for initial state $|T_{D}\rangle$ using the Fermi golden rule. (b) Dynamics at the highest resonances of (a). Solid lines represent the population in the donor sites $P_{D}$, while the x-markers illustrate the behavior of $1-P_{T_{A}}$ with $P_{T_{A}}$ being the population of $|T_{A}\rangle$. All curves correspond to the initial state $|T_{D}\rangle$. The parameters for (a) and (b) not specified in the plots are $g=\omega$, $J=0.03\omega$, $\gamma=0.015\omega$, $p=1$, and $\bar{n}=0.01$. (c) Transfer rate $k_{T}$ as a function of $\epsilon$ in the non-perturbative regime $(J\sim\lambda)$ for different initial conditions and parameter values. The parameter sets of data points marked by stars are analyzed further in panel (d). (d) Dynamics of the donor population for the parameters signaled by stars in (c). Solid lines and x-markers are defined as in (b). The parameters for (c) and (d) not specified in the plots are $g=\omega$, $J=0.3\omega$, $p=1$, and $\bar{n}=0.01$.
@@ III.1 Perturbative regime
As shown in Appendix A, this Hamiltonian corresponds to a displaced harmonic oscillator with displacement $\alpha_{j}=-g_{j}/2\omega$ (see Fig. 1). The eigenstates of the bare system are $|\psi_{j,n}\rangle=|j\rangle\otimes|n,\alpha_{j}\rangle$, where $|n,\alpha_{j}\rangle$ denote the Fock states of the displaced operator at site $j$, and the states $|j\rangle$ are defined in Eq. (4). The eigenenergies of $H_{0}$ are therefore given by:
@@
where the Franck-Condon factor $|\langle{n,\alpha_{f}}|0,\alpha_{i}\rangle|^{2}$ follows a Poissonian distribution determined by $\tilde{n}=(\alpha_{f}-\alpha_{i})^{2}=(g_{f}/2\omega-g_{i}/2\omega)^{2}$. Here, the dissipation is taken into account by considering a Lorentzian broadening of the delta functions in the original FGR [56, 55]. The total transfer rate to the acceptor site is defined by the sum over all possible acceptor states $|\psi_{f,n}\rangle$ as $k_{T}=\sum_{f\in A}k_{T}^{if}$. In this regime, the transfer is maximized when the energy is conserved, namely when $\Delta E_{if}=E_{f,n}-E_{i,0}=0$, which leads to the resonance condition $\epsilon=n\omega+(g_{f}^{2}/4\omega-g_{i}^{2}/4\omega)$.
In order to gain an intuitive understanding of the role of delocalized states in the transfer process, it is instructive to take the following choice of spin-boson coupling $g_{1}=g_{2}=-g_{3}=-g_{4}=g$. In this special case, we can include the intra-monomer nearest neighbor coupling terms in the unperturbed Hamiltonian, that is:
@@
The unperturbed eigenstates are unchanged in the vibrational degrees of freedom, but the spin eigenstates are now defined by the triplet and singlet excitonic states on the donor and acceptor sites, respectively:
@@
The new eigenstates can still be defined as $|\psi_{j,n}\rangle$ with $j=T_{D},S_{D},T_{A},S_{A}$. In this excitonic basis, the FGR in Eq. (9) considers indices $i=T_{D},S_{D}$ and $f=T_{A},S_{A}$. These eigenstates have energies identical to those on Eq.(8) with the difference that the triplet (singlet) states are shifted by $+J\;(-J)$ (see Fig. 1b). In this new notation, the perturbation Hamiltonian $H_{J}$ now corresponds to all the inter-monomer coupling terms:
@@
The coupling strengths are explicitly given by:
@@
A plot of the ratio of these inter-monomer couplings as a function of $p$ is shown in Fig. 7 in Appendix B. For values of $p\leq 3$, the coefficient $J_{TT}$ is the dominant term, with its value being significantly larger than both $J_{ST}$ and $J_{SS}$. This means that if the system is initialized in a triplet donor state $|\psi_{T_{D},0}\rangle$, the triplet-to-triplet transfer will be faster than the one associated with any initial donor product state since $J_{TT}$ is greater than any individual inter-monomer coupling. Additionally, because both triplet states, $|T_{D}\rangle$ and $|T_{A}\rangle$, have their energy shifted by $+J$, the resonance condition remains $\epsilon=n\omega$, which is independent of the spin-phonon coupling strengths $g_{i}$ for the current parameter choice.
We check this prediction numerically by integrating Eq. (2) for different initial states up until $\omega t_{\rm sim}/2\pi=100$. In Fig. 2(a-b), we show the transfer rate $k_{T}$ from donor to acceptor as a function of $\epsilon$ for $J=0.03\omega$, which is well within the perturbative regime. The transfer rate obtained numerically is defined in terms of the population in the donor state $P_{D}=\langle(|1\rangle\langle 1|+|2\rangle\langle 2|)\rangle$ as [60, 55, 44]:
@@
Fig. 2(a) shows the expected peaks in the transfer rate at the resonances $\epsilon=n\omega$ for triplet, singlet, and product states. The relative height of the peaks is determined by the Franck-Condon factor, which, following the Poissonian distribution dependent on $g$ and $\omega$ (see Eq. (9)), peaks at $\tilde{n}$=1 since $g=\omega$ for all the simulations. As expected, the numerical data shows great agreement with Eq. (9) (black solid line).
@@
In Fig. 2(b), we show the transfer dynamics from the initial state $|\psi_{T_{D},0}\rangle$ for the three main resonances $\epsilon=\omega,2\omega,3\omega$. $P_{D}$ shows an approximately exponential decay as expected from Eq. (9) into the acceptor state. Importantly, Fig. 2(b) also shows $1-P_{T_{A}}$ (signaled by x-markers), where $P_{T_{A}}=\langle|T_{A}\rangle\langle T_{A}|\rangle$, confirming that the transfer into the acceptor is a coherent process in which not only the population but the full quantum state is transferred to the acceptor site. It is worth noting that, when starting in the triplet state, the system behaves approximately as a two-level system defined by the $\{|T_{D}\rangle,|T_{A}\rangle\}$ subspace. In this electronic two-level description, an initial state $|\psi_{T_{D},0}\rangle$ is coupled to state $|\psi_{T_{A},n}\rangle$ through the electronic coupling $J_{TT}$ when $\epsilon=n\omega$. Decay through the relaxation rate $\gamma$ couples the latter to lower energy vibrational states $|\psi_{T_{A},m}\rangle$ with $m<n$. Eventually, the system decays to state $|\psi_{T_{A},0}\rangle$ where, in the case of near-zero temperature, it gets trapped as this state is off-resonant with respect to any donor state (see section  IV for the effects of finite temperature). For the population to remain trapped in $|\psi_{T_{A},0}\rangle$, it is necessary that $\epsilon>2J$, otherwise, the state $|\psi_{S_{D},0}\rangle$ could get populated even if this process is slower (rate proportional to $J_{ST}^{2}$). We adopt the condition $\epsilon>2J$ throughout this work.

### v2

# Delocalized Excitation Transfer in Open Quantum Systems with Long-Range Interactions

@@ Abstract
The interplay between coherence and system-environment interactions is at the basis of a wide range of phenomena, from quantum information processing to charge and energy transfer in molecular systems, biomolecules, and photochemical materials. In this work, we use a Frenkel exciton model with long-range interacting qubits coupled to a damped collective bosonic mode to investigate vibrationally assisted transfer processes in donor-acceptor systems featuring internal substructures analogous to light-harvesting complexes. We find that certain delocalized excitonic states maximize the transfer rate and that the entanglement is preserved during the dissipative transfer over a wide range of parameters. We investigate the reduction in transfer caused by static disorder, white noise, and finite temperature and study how transfer efficiency scales as a function of the number of dimerized monomers and the component number of each monomer, finding which excitonic states lead to optimal transfer. Finally, we provide a realistic experimental setting to realize this model in analog trapped-ion quantum simulators. Analog quantum simulation of systems comprising many and increasingly complex monomers could offer valuable insights into the design of light-harvesting materials, particularly in the non-perturbative intermediate parameter regime examined in this study, where classical simulation methods are resource-intensive.
@@ I Introduction
CT and ET in the presence of a complex environment can be generally described by Frenkel exciton models [30, 31, 32, 33] that employ discrete Hilbert spaces to encode the electronic states along with harmonic oscillators to describe molecular and solvent vibrations. Recently, it has been shown that both electronic and vibrational degrees of freedom can be directly mapped onto state-of-the-art analog quantum simulators based on trapped atomic ions [34] and superconducting qubits [35]. These platforms allow for precise control of the coherent part of the evolution [36, 37, 38, 39, 40, 41, 42] and, in some cases, also provide turning knobs to tailor the system-bath coupling implementing reservoir engineering [43, 44, 45, 46].
In this work, we theoretically and numerically investigate a minimal Frenkel exciton model, where the donor and acceptor have an internal substructure [47] similar to light-harvesting complexes in which different monomers feature strongly coupled molecular sites. Here, the molecular sites are encoded in collections of two-level systems, among which the excitation can be shared. The two-level systems are coupled to a damped collective bosonic mode, where the damping arises from the presence of an Ohmic bath.
Figure 1: Model and trapped-ion implementation: (a) A chain of 4 qubit ions and 2 coolants can realize the Frenkel exciton model (see main text). The red (blue) ions on the sides represent donor and acceptor monomers, respectively, and the information in each ion is encoded in a qubit. The two cyan ions in the middle represent the cooling ions used for the sympathetic cooling of the ion chain motion. (b) Pictorial representation of the spectrum of the total Hamiltonian in the symmetric case, described by Eqs. (10) and (12). Donor (red) and acceptor (blue) surfaces are shown as a function of the normalized reaction coordinate $y/y_{0}=1/2(a^{\dagger}+a)$ with their respective non-interacting harmonic wavefunctions. The effect of the bath is illustrated by the cyan wavy arrows that represent the relaxation with rate $\gamma$. The qubits in the donor/acceptor sites are strongly coupled and can be described in terms of local triplet and singlet states $|T_{D,A}\rangle,|S_{D,A}\rangle$ separated by an energy $2J$. These states are, in turn, coupled by inter-monomer couplings $J_{AB}$, where $AB=TT,ST,SS$ (see main text). The inter-monomer couplings $J_{AB}$ induce anti-crossings between the two surfaces allowing the donor to acceptor transfer.
@@
In addition to its fundamental theoretical interest, this model is also motivated by being experimentally realizable on an analog trapped-ion simulator, where the electronic degrees of freedom can be encoded in the internal atomic states, and the molecular vibrations in the collective vibrations of the ions around their equilibrium positions (see Section VI). Finally, in this work, we provide experimentally accessible parameters to realize this model in a trapped-ion simulator using a chain of qubit and coolant ions [44]. This paper is organized as follows. In section II, we introduce the model for excitation transfer. Section III focuses on the transfer rate of an initially delocalized excitation in a dimer, where each monomer is realized by two ions, both in the perturbative and non-perturbative regimes. Additionally, we investigate the effect of static disorder and noise on the transfer rate. In section IV, we investigate the effect of temperature on transfer rates. In section V, we discuss how this model can be extended to longer chains with more monomers and with monomers composed of more than two qubits. In section VI, we propose an experimental implementation of the model in a trapped-ion simulator. Finally, in section VII, we summarize our findings and provide an outlook for future studies.
@@ II Model
We consider a one-dimensional system of $N$ spin-$1/2$ qubit ions interleaved with $M$ coolant ions, which are used to perform reservoir engineering on the shared bosonic mode [44], as shown in Fig. 1(a). The qubit ions interact via long-range spin exchange interactions $J_{ij}$ [49] and are subject to local energy shifts $\epsilon_{i}$. At the same time, the two-level systems are coupled to the collective bosonic mode defined by creation (annihilation) operators $a^{\dagger}(a)$ with frequency $\omega$, resulting in the total Hamiltonian:
@@
The Hamiltonian $H_{\rm el}$ encodes the electronic degrees of freedom with the energy landscape that determines the roles of the molecular sites and their mutual electronic couplings. We assume a power-law decay for the electronic couplings $J_{ij}=J/d_{ij}^{p}$, that can be natively realized in trapped-ion systems with $J>0$ [49]. Here, $d_{ij}$ is the distance between the qubit ions, and $p$ is the power-law exponent. Systems in $\mathcal{D}$ dimensions with $p\leq 2\mathcal{D}+1$ have generated a lot of interest recently because deriving tight bounds on the ultimate speed of quantum information transfer is challenging in this regime [50, 51, 52, 53, 54]. Here, we consider $p=1$ for most results, which is well within this region. While the interactions with $p=1$ possess a longer range than the dipole-dipole couplings $(p=3)$, typically assumed in Frenkel exciton models, they can have physical relevance in the case of emitters separated by distances larger than the characteristic photon wavelengths [55, 56, 57]. We also comment on how our results extend to the case of $p=3$ in Appendix B. The Hamiltonian $H_{\rm v}$ describes a collective vibrational mode of motion that can be encoded in a normal mode of the ion chain, defining a reaction coordinate $y/y_{0}=1/2(a^{\dagger}+a)$, with $y_{0}=\sqrt{\hbar/2m\omega}$, and $m$ being the ion mass [58]. Finally, $H_{\rm el-v}$ represents the interaction between electron and vibration with site-dependent couplings $g_{i}$.
@@ III Two-Monomer model
Let us consider a model with only two monomers, which has been used as a minimal model of charge transfer in photosynthetic complexes [62, 63, 64, 23, 22]. In this case, each monomer is composed of two sites and separated by $M=2$ coolant ions (see section VI.2), as shown in Fig. 1. In this case, we can use the following notation:
@@
Figure 2: Excitation transfer: Transfer rate $k_{T}$ in the perturbative ((a) and (b)) and non-perturbative regimes ((c) and (d)). (a) Transfer rate $k_{T}$ as a function of $\epsilon$ in the perturbative regime ($J,\gamma\ll\omega$) for the different initial states $|T_{D}\rangle$ (T), $|S_{A}\rangle$ (S), and $|2\rangle$ (P). The black solid line represents the transfer predicted for initial state $|T_{D}\rangle$ using the Fermi Golden Rule $k_{T}^{\rm FGR}$. Note that we defined $\tilde{k}_{T}^{\rm FGR}=\frac{1}{5}k_{T}^{\rm FGR}$ where we rescale by a factor of $\frac{1}{5}$ to display it clearly in the same plot. (b) Dynamics at the highest resonances of (a). Solid lines represent the population in the donor sites $P_{D}$, while the x-markers illustrate the behavior of $1-P_{T_{A}}$ with $P_{T_{A}}$ being the population of $|T_{A}\rangle$. All curves correspond to the initial state $|T_{D}\rangle$. The parameters for (a) and (b) not specified in the plots are $g=\omega$, $J=0.03\omega$, $\gamma=0.015\omega$, $p=1$, and $\bar{n}=0.01$. (c) Transfer rate $k_{T}$ as a function of $\epsilon$ in the non-perturbative regime $(J\sim\lambda)$ for different initial conditions and parameter values. The parameter sets of data points marked by stars are analyzed further in panel (d). (d) Dynamics of the donor population for the parameters signaled by stars in (c). Solid lines and x-markers are defined as in (b). The parameters for (c) and (d) not specified in the plots are $g=\omega$, $J=0.3\omega$, $p=1$, and $\bar{n}=0.01$.
@@ III.1 Perturbative regime
As shown in Appendix A, this Hamiltonian corresponds to a displaced harmonic oscillator with displacement $\alpha_{j}=-g_{j}/2\omega$ (see Fig. 1(b)). The eigenstates of the bare system are $|\psi_{j,n}\rangle=|j\rangle\otimes|n,\alpha_{j}\rangle$, where $|n,\alpha_{j}\rangle$ denote the Fock states of the displaced operator at site $j$, and the states $|j\rangle$ are defined in Eq. (4). The eigenenergies of $H_{0}$ are therefore given by:
@@
where the Franck-Condon factor $|\langle{n,\alpha_{f}}|0,\alpha_{i}\rangle|^{2}$ follows a Poissonian distribution determined by $\tilde{n}=(\alpha_{f}-\alpha_{i})^{2}=(g_{f}/2\omega-g_{i}/2\omega)^{2}$. Here, the dissipation is taken into account by considering a Lorentzian broadening of the delta functions in the original FGR [59, 58]. The total transfer rate to the acceptor site is defined by the sum over all possible acceptor states $|\psi_{f,n}\rangle$ as $k_{T}^{\rm FGR}=\sum_{f\in A}k_{T}^{if}$. In this regime, the transfer is maximized when the energy is conserved, namely when $\Delta E_{if}=E_{f,n}-E_{i,0}=0$, which leads to the resonance condition $\epsilon=n\omega+(g_{f}^{2}/4\omega-g_{i}^{2}/4\omega)$.
In order to gain an intuitive understanding of the role of delocalized states in the transfer process, it is instructive to take the following choice of spin-boson coupling $g_{1}=g_{2}=-g_{3}=-g_{4}=g$. Although a seemingly more natural choice would have been to choose all couplings $g_{j}$ to also share the same sign, such parameter sets would, however, severely lower the transfer rates for any $\epsilon>0$, as shown in Eq. (9). This parameter choice can be implemented in a trapped-ion setup as described in Section VI. In this special parameter choice, we can include the intra-monomer nearest neighbor coupling terms in the unperturbed Hamiltonian, that is:
@@
The unperturbed vibrational eigenstates are unchanged, but the spin eigenstates are now defined by the triplet and singlet excitonic states on the donor and acceptor sites, respectively:
@@
The new eigenstates can still be defined as $|\psi_{j,n}\rangle$ with $j=T_{D},S_{D},T_{A},S_{A}$. In this excitonic basis, the FGR in Eq. (9) considers indices $i=T_{D},S_{D}$ and $f=T_{A},S_{A}$. These eigenstates have energies identical to those on Eq.(8) with the difference that the triplet (singlet) states are shifted by $+J\;(-J)$. A pictorial representation of the energetics in this new basis is provided in Fig. 1(b). In this new notation, the perturbation Hamiltonian $H_{J}$ now corresponds to all the inter-monomer coupling terms (see green arrows in Fig. 1(b)):
@@
We note that there are no intra-monomer couplings in this basis, thanks to the values chosen for $g_{j}$. The coupling strengths are explicitly given by:
@@
A plot of the ratio of these inter-monomer couplings as a function of $p$ is shown in Fig. 7 in Appendix B. For values of $p\leq 3$, the coefficient $J_{TT}$ is the dominant term, with its value being significantly larger than both $J_{ST}$ and $J_{SS}$. This means that if the system is initialized in a triplet donor state $|\psi_{T_{D},0}\rangle$, the triplet-to-triplet transfer will be faster than the one associated with any initial donor product state since $J_{TT}$ is greater than any individual inter-monomer coupling. Additionally, because both triplet states, $|T_{D}\rangle$ and $|T_{A}\rangle$, have their energy shifted by $+J$, the resonance condition remains $\epsilon=n\omega$, which is independent of the spin-phonon coupling strengths $g_{j}$ for the current parameter choice (see Fig. 1(b)).
We check this prediction numerically by integrating Eq. (2) for different initial states up until $\omega t_{\rm sim}/2\pi=100$ (see Appendix C for details on the dynamics of the bosonic degrees of freedom and truncation of the bosonic Hilbert space). In Fig. 2(a-b), we show the transfer rate $k_{T}$ from donor to acceptor as a function of $\epsilon$ for $J=0.03\omega$, which is well within the perturbative regime. The transfer rate obtained numerically is defined in terms of the population in the donor state $P_{D}=\langle(|1\rangle\langle 1|+|2\rangle\langle 2|)\rangle$ as [65, 58, 44]:
@@
Fig. 2(a) shows the expected peaks in the transfer rate at the resonances $\epsilon=n\omega$ for triplet, singlet, and product states. The relative height of the peaks is determined by the Franck-Condon factor, which, following the Poissonian distribution dependent on $g$ and $\omega$ (see Eq. (9)), peaks at $\tilde{n}$=1 since $g=\omega$ for all the simulations. We note that we choose this value of $g$ to explore the regime in which the donor and acceptor sites are described by displaced potential energy surfaces with quantized vibrational levels (see Fig. 1(b)). The weakly coupled regime with $g/\omega\lesssim 0.1$ has been explored before, see, for example, Refs.[36, 66]. The numerically obtained transfer rates qualitatively agree with those predicted by the FGR, as shown by the black solid line (note that we display $\tilde{k}_{T}^{\rm FGR}=k_{T}^{\rm FGR}/5$ in the plot). However, the FGR predicts a higher transfer rate, which can be attributed to the fact that this expression is derived under the assumption $\gamma\gg J$ [65]. This condition differs from the parameters used in panel (a), where $J=0.03\omega$ and $\gamma=0.015\omega$.
@@
In Fig. 2(b), we show the transfer dynamics from the initial state $|\psi_{T_{D},0}\rangle$ for the three main resonances $\epsilon=\omega,2\omega,3\omega$. $P_{D}$ shows an approximately exponential decay as expected from Eq. (9) into the acceptor state. Importantly, Fig. 2(b) also shows $1-P_{T_{A}}$ (signaled by x-markers), where $P_{T_{A}}=\langle|T_{A}\rangle\langle T_{A}|\rangle$, confirming that the transfer into the acceptor is a coherent process in which not only the population but the full quantum state is transferred to the acceptor site. Given that $|T_{D}\rangle$ and $|T_{A}\rangle$ are maximally entangled states of the qubits in the donor and acceptor, respectively, the dynamics can also be understood in terms of entanglement transfer, see Appendix D for more details. It is worth noting that, when starting in the triplet state, the system behaves approximately as a two-level system defined by the $\{|T_{D}\rangle,|T_{A}\rangle\}$ subspace. In this electronic two-level description, an initial state $|\psi_{T_{D},0}\rangle$ is coupled to state $|\psi_{T_{A},n}\rangle$ through the electronic coupling $J_{TT}$ when $\epsilon=n\omega$. Decay through the relaxation rate $\gamma$ couples the latter to lower energy vibrational states $|\psi_{T_{A},m}\rangle$ with $m<n$. Eventually, the system decays to state $|\psi_{T_{A},0}\rangle$ where, in the case of near-zero temperature, it gets trapped as this state is off-resonant with respect to any donor state (see section  IV for the effects of finite temperature). For the population to remain trapped in $|\psi_{T_{A},0}\rangle$, it is necessary that $\epsilon>2J$, otherwise, the state $|\psi_{S_{D},0}\rangle$ could get populated even if this process is slower (rate proportional to $J_{ST}^{2}$). We adopt the condition $\epsilon>2J$ throughout this work.